# Boxed in from all sides: a global fit to loopy dark matter and neutrino masses

### v1

@@ I Introduction
This provides exciting opportunities for understanding the underlying properties of both dark matter and neutrinos. However, it presents a challenge to probe given the feeble interactions of both sectors with the rest of the SM. Still, significant work has gone into investigating such a possibility. A large coupling between dark matter and neutrinos can leave imprints on cosmological observables such as the CMB and large scale structure Boehm et al. (2001); Boehm and Schaeffer (2005); Mangano et al. (2006); Serra et al. (2009); van den Aarssen et al. (2012); Wilkinson et al. (2014); Schewtschenko et al. (2015); Boehm et al. (2014); Escudero et al. (2015); Olivares-Del Campo et al. (2018); Balducci et al. (2018); Wilkinson et al. (2014); Bertoni et al. (2015); Di Valentino et al. (2018); Diacoumis and Wong (2019); Mangano et al. (2006); Serra et al. (2010); Hooper and Lucca (2022). It may also manifest itself in galaxies Akita and Ando (2023); Cline and Puel (2023) as well as other astrophysical processes. For instance, ultra high energy neutrinos of astrophysical origin can scatter with ambient dark matter and lead to attenuation of the expected neutrino spectrum at Earth Reynoso and Sampayo (2016); Argüelles et al. (2017); Vincent et al. (2017); Argüelles et al. (2020); Choi et al. (2019). Moreover, lower energy neutrinos, originating from stars, supernovae or the cosmic neutrino background can scatter with ambient dark matter, resulting in both a modification of the expected spectrum at Earth Farzan and Palomares-Ruiz (2014, 2019); Carpio et al. (2023); Balantekin et al. (2023) as well as up-scattering of dark matter Zhang (2022); Das and Sen (2021); Jho et al. (2021); Ghosh et al. (2022). These searches are complemented by model-independent investigations of galactic and extra-galactic dark matter annihilating or decaying directly into neutrinos (see refs. Argüelles et al. (2021); Buckley et al. (2023); Argüelles et al. (2023); Das et al. (2023); Fiorillo et al. (2023); Song et al. (2024); Allahverdi et al. (2024); Fuß et al. (2024); Das et al. (2024) and references therein).
@@
On the theory side, the landscape of scenarios that provide an explanation for dark matter-neutrino interactions is broad. Many of the underlying models provide a viable candidate for dark matter while explaining the origin of neutrino masses, and provide complementary signals in terrestrial experiments. Some common examples include the scotogenic and scotogenic-like models (see ref. Ma (1998, 2006a); Boehm et al. (2008) and references thereafter), and the Zee model (see ref. Zee (1980) and references thereafter). These scenarios are well-motivated, but come with many parameters, making them hard to constrain. It is common in the literature to fix some parameters to ad hoc benchmark values while varying others to show the viable parameter space. This can lead to misinterpretations of the parameter space, depending on the parameters represented and their constraints.
@@
To explore the rich phenomenology presented by this model, we perform a global likelihood analysis and investigate the terrestrial and astrophysical constraints on the multi-dimensional parameter space, without fixing any parameters. Remarkably, we will find that for couplings that are small enough for the theory to remain perturbative, constraints from cosmology, indirect detection, neutrino oscillation and collider experiments bound this model to a closed region of parameter space. This model naturally points to an MeV–GeV mass dark matter candidate $\eta$.
@@ II Model
where $\lambda^{\prime}$ and $\lambda$ are the quartic couplings linking the DM to the complex scalar mediator, $Y_{\nu}$ is a Yukawa-type coupling linking the dark sector to the left-handed neutrinos of the SM. $\mu$ is a constant that splits the U(1) symmetry of $\Phi$ as in the SLIM mechanism of Ref. Boehm et al. (2008), ultimately leading to the neutrino mass terms. The quartic coupling $\lambda^{\prime}$ is neglected for simplicity as in Ref. Chao (2020).
@@
The quantities $m^{2}_{\phi_{1,2}}=m_{\Phi}^{2}\pm\mu^{2}$ are the masses of $\phi_{1,2}$ respectively, and by construction, $m_{\phi_{1}}>m_{\phi_{2}}$. This phenomenology can be straightforwardly generalised to three flavours. However, for simplicity, we focus on a single mass eigenstate at a time.
@@
The direct implication is that neutrino masses are generated by a Lepton Number Violating (LNV) process. Setting $m_{\phi_{1}}=m_{\phi_{2}}$ restores lepton number conservation and neutrino masses vanish. Going forward, we parametrise this with the dimensionless quantity
@@ III.4 Dark matter-neutrino elastic scattering
Many recent references have revisited these constraints, but mainly focus on cross sections that are constant with $T$. As we will find in Sec. IV, these elastic scattering constraints will always be subdominant to the others discussed here, so they will not be explored further.
@@ IV Results: Global analysis
Figure 6: 95.45% C.R. posterior constraints from the four MCMC parameter scans. Top: couplings-mediator mass plane; bottom: mass splitting–dark matter mass. Line styles represent coupling to the heaviest neutrino eigenstates ($\nu_{3}$, dashed) and lightest ($\nu_{3}$, solid), assuming normal ordering. The light color shows dark matter $\eta$ that saturates the DM relic abundance, while the dark color is allowed to be underabundant.
@@
These constraints are shown more explicitly in Figs. 7, corresponding to the $\lambda Y^{2}_{\nu}$$m_{\phi_{1}}+m_{N}$ plane, and 8, showing the $\delta m_{\phi}-m_{\eta}$ plane. In these figures, we show the 95.45% confidence level (frequentist) profile likelihood ratio contour as a solid line, and the 95.45% (bayesian) credible region as a dashed line. The two generally trace each other well.
@@
Fig. 7 shows the plane of the couplings versus mediator masses for the four different models, and the color scale indicates the dark matter mass $m_{\eta}$ for each of the samples. The parameter space is clearly bounded from all sides by the constraints, as described above, represented by shaded regions. These are overlaid to indicate the dominant constraint in each region. In all cases, the couplings and mediator masses are bounded from below by a combination of the relic abundance, BBN, annihilation cross-section and kaon decay. As the mediator masses and couplings increase, larger mass splittings and lower dark matter masses are needed to keep dark matter from being overproduced. The low DM masses in turn are constrained in a small range by the Super-K limit on the annihilation cross-section and below by BBN.
The Saturated DM + Heavy $\nu$ (top left) and Saturated DM + Light $\nu$ (bottom left) are bounded from above by the perturbativity of the couplings in combination with the DM relic abundance. The bounds are slightly different for both cases due to limits on the mass of the heaviest and lightest neutrino, respectively. The Underabundant DM + Heavy $\nu$ (top right) and Underabundant DM + Light $\nu$ (bottom right) are bounded from above by the perturbativity of the couplings in combination with the limits on the Z decay to $N\phi_{1}\nu$.
@@
In Fig. 8 all panels show that the model naturally favors lower ($\sim 10$ MeV-GeV) mass dark matter in order to generate non-zero neutrino masses and/or be consistent with relic abundance and perturbativity constraints. Furthermore, the mass splitting is bounded from above by the combination of BBN and the upper limit of the neutrino masses. In the Saturated DM + Heavy $\nu$ (top left) and Underabundant DM + Heavy $\nu$ (top right) cases the mass splitting is bounded from below by the lower limit on the mass of the heaviest neutrino and the perturbativity of the couplings.
In these cases, we notice that the mass splitting is roughly inversely proportional to the scalar mediator mass (as seen in the pink color scale), so smaller mass splittings lead to larger mediator masses in order to reproduce the mass of the heaviest neutrino. Additionally, larger couplings are needed which leads to the perturbativity limit. In the Saturated DM case, the upper limit on the DM mass is tighter since it is not allowed to be underproduced.
@@ V Discussion and Conclusion
We have performed a detailed phenomenological analysis of a model in which scalar dark matter interacts with neutrinos, primarily at next-to-leading order in perturbation theory. This scenario self-consistently provides an explanation for light thermal relic dark matter as well as a neutrino mass generation mechanism. We have shown that various cosmological, astrophysical and experimental observables bound this model, on all sides, leaving only a finite volume of parameter space available. We find that the dark matter relic abundance and perturbativity constraints naturally lead to a sub-GeV dark matter candidate, constraining the model from the right. At the same time, BBN prevents the dark matter from being lighter than $\sim 10$ MeV, while heavy meson decays require the sum of the mediator masses to be heavier than $\sim 500$ MeV, bounding the parameter space from the left. Finally, an upper limit on the neutrino mass and rare Z-boson decays bound the space from above.
The bounds on this dark sector coupling to the lightest and heaviest neutrinos are fairly robust, and the constraints on the parameter space are only relaxed by a small amount when $\eta$ is allowed to be underabundant. Taken together, these observations hint that a full, 3-flavor model should not be so different. Such an analysis would come at the cost of many more free parameters and correlations, but the generalization is in principle straightforward. Additional constraints (on the 3-flavor model)—and opportunities—would of course arise from neutrino oscillation, these are beyond the scope of this work.
A remaining question, then, is how detectable such an extension of the standard model is in the near future. A more precise measurement of the relic abundance $\Omega_{DM}h^{2}$ will of course narrow the range of couplings and mediator masses, while next generation oscillation parameter measurements will help pin down the mass splittings. Upcoming experiments, such as DUNE Buckley et al. (2023), JUNO Argüelles et al. (2021) and Hyper-Kamiokande Bell et al. (2020), could offer improvements by providing sensitivity to dark matter-neutrino annihilation cross-sections; a neutrino line signal would indeed be a smoking gun for a neutrinophillic dark sector.
@@ Acknowledgements
We thank Richard Ruiz, Cristina Mondino, Matheus Hostert, Vedran Brdar, Martín de los Rios and the TheDeAs group for useful discussions. All authors acknowledge support from the Arthur B. McDonald Canadian Astroparticle Physics Research Institute. KMC is supported by “la Caixa” Foundation through the fellowship LCF/BQ/DI22/11940024 and by the Spanish Agencia Estatal de Investigación through the grant PID2021-125331NB- I00. During completion of this work, GM was supported in part by the UC office of the President through the UCI Chancellor’s Advanced Postdoctoral Fellowship, the U.S. National Science Foundation under Grant PHY-2210283 and the Natural Sciences and Engineering Research Council of Canada (NSERC). ACV is supported by NSERC and the province of Ontario through an ERA, with equipment funded by the Canada Foundation for Innovation and the Province of Ontario, and housed at the Queen’s Centre for Advanced Computing. Research at Perimeter Institute is supported by the Government of Canada through the Department of Innovation,Science, and Economic Development, and by the Province of Ontario.
@@ A.1 Dark matter-neutrino elastic scattering
In the frame of the dark matter, that is basically stationary, the differential cross-section is Argüelles et al. (2017):
@@
is the Fermi-Dirac distribution for the neutrinos.
Numerically, the thermally-averaged elastic scattering cross-section is calculated as:
@@ Appendix B Z boson decay width
$\text{Integrand}(m_{12}^{2},\alpha,\beta,Y_{\nu})=$ $\frac{\alpha_{l}\pi Y_{\nu}^{2}}{3\cos^{2}\theta_{w}m_{Z}^{2}\sin^{2}\theta_{w}}(\cos^{2}\theta_{w}+\sin^{2}\theta_{w})^{2}\times$ (44)
@@
All the code to reproduce the results presented in this work is available at https://github.com/Karen-Macias/DM-neutrino-loop.

### v2

@@ I Introduction
This provides exciting opportunities for understanding the underlying properties of both dark matter and neutrinos. However, it presents a challenge to probe given the feeble interactions of both sectors with the rest of the SM. Still, significant work has gone into investigating such a possibility. A large coupling between dark matter and neutrinos can leave imprints on cosmological observables such as the CMB and large scale structure Boehm et al. (2001); Boehm and Schaeffer (2005); Mangano et al. (2006); Serra et al. (2010a); van den Aarssen et al. (2012); Wilkinson et al. (2014); Schewtschenko et al. (2015); Boehm et al. (2014); Escudero et al. (2015); Olivares-Del Campo et al. (2018); Balducci et al. (2018); Wilkinson et al. (2014); Bertoni et al. (2015); Di Valentino et al. (2018); Diacoumis and Wong (2019); Mangano et al. (2006); Serra et al. (2010b); Hooper and Lucca (2022). It may also manifest itself in galaxies Akita and Ando (2023); Cline and Puel (2023) as well as other astrophysical processes. For instance, ultra high energy neutrinos of astrophysical origin can scatter with ambient dark matter and lead to attenuation of the expected neutrino spectrum at Earth Reynoso and Sampayo (2016); Argüelles et al. (2017a); Vincent et al. (2017); Argüelles et al. (2020); Choi et al. (2019). Moreover, lower energy neutrinos, originating from stars, supernovae or the cosmic neutrino background can scatter with ambient dark matter, resulting in both a modification of the expected spectrum at Earth Farzan and Palomares-Ruiz (2014, 2019); Carpio et al. (2023); Balantekin et al. (2023) as well as up-scattering of dark matter Zhang (2022); Das and Sen (2021); Jho et al. (2021); Ghosh et al. (2022). These searches are complemented by model-independent investigations of galactic and extra-galactic dark matter annihilating or decaying directly into neutrinos (see refs. Argüelles et al. (2021); Buckley et al. (2023); Argüelles et al. (2023); Das et al. (2023); Fiorillo et al. (2023); Song et al. (2024); Allahverdi et al. (2024); Fuß et al. (2024); Das et al. (2024) and references therein).
@@
On the theory side, the landscape of scenarios that provide an explanation for dark matter-neutrino interactions is broad. Many of the underlying models provide a viable candidate for dark matter while explaining the origin of neutrino masses, and provide complementary signals in terrestrial experiments. Some common examples include the scotogenic and scotogenic-like models (see ref. Ma (1998, 2006a); Boehm et al. (2008), more recently Herms et al. (2023); Jana et al. (2024), and Borah et al. (2019) for a connection to leptogenesis), and the Zee model (see ref. Zee (1980) and references thereafter). These scenarios are well-motivated, but come with many parameters, making them hard to constrain. It is common in the literature to fix some parameters to ad hoc benchmark values while varying others to show the viable parameter space. This can lead to misinterpretations of the parameter space, depending on the parameters represented and their constraints.
@@
To explore the rich phenomenology presented by this model, we perform a global likelihood analysis and investigate the terrestrial and astrophysical constraints on the multi-dimensional parameter space, without fixing any parameters. Remarkably, we will find that for couplings that are small enough for the theory to remain perturbative, constraints from cosmology, indirect detection, neutrino oscillation, and collider experiments bound this model to a closed region of parameter space. This model naturally points to an MeV–GeV mass dark matter candidate $\eta$. While this model may contain a second region of parameter space at lower couplings leading to dark matter production via freeze-in (as opposed to freeze-out), this would likely obviate some of the benefits mentioned above. We leave this technically distinct exploration for future work.
@@ II Model
where $\lambda^{\prime}$ and $\lambda$ are the quartic couplings linking the DM to the complex scalar mediator, $Y_{\nu}$ is a Yukawa-type coupling linking the dark sector to the left-handed neutrinos of the SM. $\mu$ is a constant that splits the U(1) symmetry of $\Phi$ as in the SLIM mechanism of Ref. Boehm et al. (2008), ultimately leading to the neutrino mass terms. Interactions between the SM Higgs and the new scalars, as well as the quartic coupling $\lambda^{\prime}$, are neglected for simplicity as in Ref. Chao (2020).
@@
The quantities $m^{2}_{\phi_{1,2}}=m_{\Phi}^{2}\pm\mu^{2}$ are the masses of $\phi_{1,2}$ respectively, and by construction, $m_{\phi_{1}}>m_{\phi_{2}}$. This phenomenology can be straightforwardly generalized to three flavors. However, for simplicity, we focus on a single mass eigenstate at a time.
@@
The direct implication is that neutrino masses are generated by a Lepton Number Violating (LNV) process. Setting $m_{\phi_{1}}=m_{\phi_{2}}$ restores lepton number conservation, and neutrino masses vanish. Going forward, we parametrize this with the dimensionless quantity
@@ III.3 Primordial Nucleosynthesis
As mentioned in the introduction, scotogenic models have been linked to leptogenesis. Our effective low-energy lagrangian Eq. (2) does not directly couple to the charged leptons, so a full baryogenesis and leptogenesis analysis is not possible. Models of leptogenesis can produce a large asymmetry in the neutrino sector Simha and Steigman (2008). We do not include such a constraint, since, again, it would require a more UV-complete model, and including all three neutrino flavors simultaneously.
@@ III.4 Dark matter-neutrino elastic scattering
Many recent references have revisited these constraints, but mainly focus on cross sections that are constant with $T$ Ferrer et al. (2023); Cline et al. (2023); Fujiwara and Herrera (2024). As we will find in Sec. IV, these elastic scattering constraints will always be subdominant to the others discussed here, so they will not be explored further.
@@ IV Results: Global analysis
Figure 6: 95.45% C.R. posterior constraints from the four MCMC parameter scans. Top: couplings-mediator mass plane; bottom: mass splitting–dark matter mass. Line styles represent coupling to the heaviest neutrino eigenstates ($\nu_{3}$, dashed) and lightest ($\nu_{1}$, solid), assuming normal ordering. The light color shows dark matter $\eta$ that saturates the DM relic abundance, while the dark color is allowed to be underabundant.
@@
These constraints are shown more explicitly in Figs. 7, corresponding to the $\lambda Y^{2}_{\nu}-m_{\phi_{1}}+m_{N}$ plane, and 8, showing the $\delta m_{\phi}-m_{\eta}$ plane. In these figures, we show the 95.45% confidence level (frequentist) profile likelihood ratio contour as a solid line, and the 95.45% (bayesian) credible region as a dashed line. The two generally trace each other well.
@@
Fig. 7 shows the plane of the couplings versus mediator masses for the four different cases, and the color scale indicates the dark matter mass $m_{\eta}$ for each of the samples. The parameter space is clearly bounded from all sides by the constraints, as described above, represented by shaded regions. These are overlaid to indicate the dominant constraint in each region. In all cases, the couplings and mediator masses are bounded from below by a combination of the relic abundance, BBN, annihilation cross-section and kaon decay. As the mediator masses and couplings increase, larger mass splittings and lower dark matter masses are needed to keep dark matter from being overproduced. The low DM masses in turn are constrained in a small range by the Super-K limit on the annihilation cross-section and below by BBN.
We choose to focus on a WIMP-like model where dark matter is produced thermally through a freeze-out mechanism. Therefore, it is natural for bigger couplings to be preferred, but the parameter space is closed at $m_{\phi_{1}}+m_{N}<500$ MeV by the constraint from kaon decay. As mentioned in the introduction, we opt for concentrating on this region and leave the question of other dark matter production mechanisms like freeze-in for future work.
The Saturated DM + Heavy $\nu$ (top left) and Saturated DM + Light $\nu$ (bottom left) cases are bounded from above by the perturbativity of the couplings in combination with the DM relic abundance. The bounds are slightly different for both cases due to limits on the mass of the heaviest and lightest neutrino, respectively. The Underabundant DM + Heavy $\nu$ (top right) and Underabundant DM + Light $\nu$ (bottom right) cases are bounded from above by the perturbativity of the couplings in combination with the limits on the Z decay to $N\phi_{1}\nu$.
@@
In Fig. 8 all panels show that the model naturally favors lower ($\sim 10$ MeV-GeV) mass dark matter in order to generate non-zero neutrino masses and/or be consistent with relic abundance and perturbativity constraints. Furthermore, the mass splitting is bounded from above by the combination of BBN and the upper limit of the neutrino masses. In the Saturated DM + Heavy $\nu$ (top left) and Underabundant DM + Heavy $\nu$ (top right) cases, the mass splitting is bounded from below by the lower limit on the mass of the heaviest neutrino and the perturbativity of the couplings.
In these cases, we notice that the mass splitting is roughly inversely proportional to the scalar mediator mass (as seen in the pink color scale), so smaller mass splittings lead to larger mediator masses in order to reproduce the mass of the heaviest neutrino. Additionally, larger couplings are needed, which leads to the perturbativity limit. In the Saturated DM case, the upper limit on the DM mass is tighter since it is not allowed to be underproduced.
@@ V Discussion and Conclusion
We have performed a detailed phenomenological analysis of a model in which scalar dark matter interacts with neutrinos, primarily at next-to-leading order in perturbation theory. This scenario self-consistently provides an explanation for light thermal relic dark matter as well as a neutrino mass generation mechanism. We have shown that various cosmological, astrophysical, and experimental observables bound this model, on all sides, leaving only a finite volume of parameter space available. We find that the dark matter relic abundance and perturbativity constraints naturally lead to a sub-GeV dark matter candidate, constraining the model from the right. At the same time, BBN prevents the dark matter from being lighter than $\sim 10$ MeV, while heavy meson decays require the sum of the mediator masses to be heavier than $\sim 500$ MeV, bounding the parameter space from the left. Finally, an upper limit on the neutrino mass and rare Z-boson decays bound the space from above.
The bounds on this dark sector coupling to the lightest and heaviest neutrinos are fairly robust, and the constraints on the parameter space are only relaxed by a small amount when $\eta$ is allowed to be underabundant. Taken together, these observations hint that a full 3-flavor model should not be so different. Such an analysis would come at the cost of many more free parameters and correlations, but the generalization is in principle straightforward. Additional constraints (on the 3-flavor model)—and opportunities—would of course arise from neutrino oscillation, these are beyond the scope of this work. Similarly, a more UV-complete model could potentially tie in with leptogenesis, unifying three major questions in particle cosmology.
A remaining question, then, is how detectable such an extension of the standard model is in the near future. A more precise measurement of the relic abundance $\Omega_{DM}h^{2}$ will of course narrow the range of couplings and mediator masses, while next-generation oscillation parameter measurements will help pin down the mass splittings. Upcoming experiments, such as DUNE Buckley et al. (2023), JUNO Argüelles et al. (2021) and Hyper-Kamiokande Bell et al. (2020), could offer improvements by providing sensitivity to dark matter-neutrino annihilation cross-sections; a neutrino line signal would indeed be a smoking gun for a neutrinophillic dark sector.
@@ Acknowledgements
We thank Richard Ruiz, Cristina Mondino, Matheus Hostert, Yasaman Farzan, Vedran Brdar, Martín de los Rios and the TheDeAs group for useful discussions. All authors acknowledge support from the Arthur B. McDonald Canadian Astroparticle Physics Research Institute. KMC is supported by “la Caixa” Foundation through the fellowship LCF/BQ/DI22/11940024 and by the Spanish Agencia Estatal de Investigación through the grants PID2021-125331NB-I00 and CEX2020-001007-S. During completion of this work, GM was supported in part by the UC office of the President through the UCI Chancellor’s Advanced Postdoctoral Fellowship, the U.S. National Science Foundation under Grant PHY-2210283 and the Natural Sciences and Engineering Research Council of Canada (NSERC). ACV is supported by NSERC and the province of Ontario through an ERA, with equipment funded by the Canada Foundation for Innovation and the Province of Ontario, and housed at the Queen’s Centre for Advanced Computing. Research at Perimeter Institute is supported by the Government of Canada through the Department of Innovation,Science, and Economic Development, and by the Province of Ontario.
@@ A.1 Dark matter-neutrino elastic scattering
In the frame of the dark matter, that is basically stationary, the differential cross-section is Argüelles et al. (2017b):
@@
is the Fermi-Dirac distribution for the neutrinos. Numerically, the thermally-averaged elastic scattering cross-section is calculated as:
@@ Appendix B Z boson decay width
Integrand $(m_{12}^{2},\alpha,\beta,Y_{\nu})=\frac{\alpha_{l}\pi Y_{\nu}^{2}}{3\cos^{2}\theta_{w}m_{Z}^{2}\sin^{2}\theta_{w}}(\cos^{2}\theta_{w}+\sin^{2}\theta_{w})^{2}\times$ (44)
@@
To reproduce the results presented in this work visit Macías Cárdenas .